# Low-Acceleration Gravitational Anomaly from Bayesian 3D Modeling of Wide Binary Orbits: Methodology and Results with Gaia DR3

### v1

@@ Abstract
Isolated wide binary stars provide natural laboratories to directly test or measure weak gravity for Newtonian acceleration $g_{\rm{N}}\lesssim 10^{-9}$ m s-2. Recent statistical analyses of wide binaries have been performed only with sky-projected relative velocities $v_{p}$ in the pairs. A new method of Bayesian orbit modeling exploiting three relative velocity components including the radial (line-of-sight) component $v_{r}$ is developed to measure a gravitational anomaly parameter $\Gamma\equiv\log_{10}\sqrt{G_{\rm{eff}}/G_{\rm{N}}}$ where $G_{\rm{eff}}$ is the effective gravitational constant for pseudo-Newtonian elliptical orbits, while $G_{\rm{N}}$ is Newton’s constant. The method infers individual probability distributions of $\Gamma$ and then combines the independent distributions to obtain a consolidated distribution in a specific range of $g_{\rm{N}}$. Here the method is described and applied to a sample of 312 wide binaries in a broad dynamic range $10^{-11.0}\lesssim g_{\rm{N}}\lesssim 10^{-6.7}$ m s-2 with $v_{r}$ uncertainties in the range $380168<\sigma_{v_{r}}<$ 380 m s-1 selected from the Gaia DR3 database. The following results are obtained: $\Gamma=0.000\pm 0.011$ ($N_{\rm{binary}}=125$) for a high acceleration regime ($10^{-7.9}\lesssim g_{\rm{N}}\lesssim 10^{-6.7}$ m s-2) agreeing well with Newton, but $\Gamma=0.085\pm 0.040$ (35) for a MOND regime ($10^{-11.0}\lesssim g_{\rm{N}}\lesssim 10^{-9.5}$ m s-2) and $\Gamma=0.063\pm 0.015$ (111) for a MOND+transition regime ($10^{-11.0}\lesssim g_{\rm{N}}\lesssim 10^{-8.5}$ m s-2). These results show that gravitational anomaly is evident for $g_{\rm{N}}\lesssim 10^{-9}$ m s-2 and $\Gamma$ in the MOND regime ($\lesssim 10^{-9.5}$ m s-2) agrees with the first-tier prediction ($\approx 0.07$) of MOND gravity theories. Finally, simulations indicate that $\Gamma$ in the MOND regime can be measured more accurately if radial velocities are better measured in the future.
@@ 1 Introduction
The nature of gravity in the low acceleration limit of the nonrelativistic regime is of great importance because any deviation from standard gravity of Newton and Einstein represented by Poisson’s equation will have profound implications for the concept of dark matter, dynamics of astrophysical systems, cosmological theories, and fundamental theories of physics. For the past decade or so, wide binary stars have been recognized (Hernandez et al., 2012; Pittordis & Sutherland, 2018; Banik & Zhao, 2018) as potentially decisive testbeds of gravity at low acceleration because a significant number of wide binaries can be found in the solar neighborhoods, permitting relatively direct observational studies and their internal dynamics is free from the effects of hypothetical dark matter.
Several statistical analyses of wide binary internal motions carried out during 2023 - 2024 based on Gaia data release 3 (DR3; Vallenari et al., 2023), by Chae (Chae, 2023a, 2024a, 2024c) and independently by Hernandez et al. (Hernandez, 2023; Hernandez et al., 2024a), indicate that standard gravity is not only broken at low internal Newtonian acceleration ($g_{\rm{N}}\lesssim 10^{-9}$ m s-2) but gravity is boosted by approximately 40% at $g_{\rm{N}}\lesssim 10^{-10}$ m s-2 in agreement with the generic prediction of modified gravity theories (Bekenstein & Milgrom, 1984; Milgrom, 2010) in the paradigm of modified Newtonian dynamics (MOND; Milgrom, 1983). Although divergent results (Banik et al., 2024) were published during the same period, the recent critical review by Hernandez et al. (2024b) (see also Chae 2024c for more details) addresses the origin of the divergent results, in particular the lack of the self-calibration procedure for the fraction of hierarchical systems in an impure binary sample (i.e. including hierarchical systems).
@@
This paper is structured as follows. In Section 2, I describe the methodology and data. In detail, I first develop a Bayesian methodology to infer probability density functions (PDFs) of gravity from individual binaries (Section 2.1), and then describe the method of statistically combining individual PDFs to obtain a consolidated probability distribution by maximizing information (Section 2.2). Section 2.3 describes the selection of a pure binary sample from the Gaia DR3 database that will be used for individual Bayesian modeling. The selection of the sample will be more stringent than previous samples that were used mainly for median statistics in bins of accelerations (or equivalent variables) by Chae (2024a, c). In Section 3, I describe the results of Bayesian modeling based on the Gaia sample. In Section 3.1, I first examine the individual Bayesian results and the overall properties of the results. In Section 3.2, I present the consolidated probability distributions of gravity with wide binaries in a Newtonian (relatively strong internal acceleration) regime. In Section 3.3, I present the main results of this work, i.e. the consolidated probability distributions of gravity in low-acceleration regimes. In Section 4, I discuss the meaning of the results and future wide binary gravity research. In Section 5, I present the current conclusion and prospects with future data on radial velocities. In Appendix A, a general transformation with the Euler angles between the orbital plane and the observer’s frame is described as it is needed to produce mock samples.
For a quick grasp of this rather lengthy paper, the reader may read Sections 2.1, 2.2, 3.2, 3.3, and 5. The wide binary samples, the Bayesian modeling results, and key Python codes will be available at a Zenodo site.
@@ 2.1 Bayesian 3D modeling
which is related to the inclination angle $i$ by $\tau=\tan i$. Finally, the magnitude of the relative physical velocity in the 3D space is then given by
@@
In Equations (10), (12), and (13), $e$ is the eccentricity, $i$ is an inclination parameter, [^1] I note that this parameter is not the same as the angle between the orbital plane and the plane of the sky. $\phi_{0}$ is the longitude of the periastron, and $\phi$ is the phase angle given by $\phi=\tan^{-1}((\Delta y^{\prime}/\Delta x^{\prime})/\cos i)$. In Equation (10), $\gamma_{g}$ is the gravity boost factor defined in Equation (1). In Equation (11), $M_{\rm{tot}}$ is the total mass of the binary system to be constrained, while $M_{\rm{tot,obs}}$ is the observational estimate based on the measured luminosities of the stars and the mass-luminosity relation given in Figure 7 of Chae (2023a) (the standard choice). The set of free parameters for each binary is thus given by $\mathbf{\Theta}=\{e,\phi_{0},i,\log_{10}(f_{M}),\Gamma\}$.
@@
For eccentricity $e$ I consider the range $0.9990.001<e<$ 0.999 with either a flat prior or a “thermal” probability density $p(e)=2e$. If the data can determine the orbit well, the flat prior may be sufficient. However, with snapshot data, $e$ is ill-constrained in particular for relatively large uncertainties of $v_{r}$ as in many cases based on the current Gaia DR3 database. When $e$ is ill-constrained (or essentially unconstrained in some cases), the prior matters because the flat prior would be equivalent to assuming that the median eccentricity is $0.5$ while the thermal prior has a median of $\sqrt{0.5}$ more consistent with the observed statistical properties. For this reason, the thermal prior will be the default choice. Note, however, that according to a Bayesian inference of eccentricities for Gaia DR3 wide binaries based on the sky-projected velocities and displacement vectors, Hwang et al. (2022) obtain somewhat higher eccentricities than the thermal distribution. Because higher eccentricities are in favor of MOND rather than Newton, this work will be slightly biased towards Newton (or a lower value of $\Gamma$), although this difference is negligible compared to the difference between the flat and thermal distributions.
@@
where $\phi=\tan^{-1}[(\Delta y^{\prime}/\Delta x^{\prime})/\cos i]$. The parameter $\phi_{0}$ is allowed to vary freely within the range $(-360^{\circ},360^{\circ})$.
For $f_{M}$, I assume a Gaussian probability density for $\log_{10}(f_{M})$ with $(\mu,\sigma)=(0,0.021)$ allowing a 5% scatter of $M_{\rm{tot}}$, which covers the alternative mass-luminosity relations shown in Figure 7 of Chae (2023a). For $i$ and $\Gamma$, I assume flat priors in the ranges: $-90^{\circ}<i<90^{\circ}$ and $-1<\Gamma<1$.
@@
Given that the averaged PDF is not satisfactory, what would be the best way to extract the maximal information from the set of individual PDFs, or how to combine independent measurements of the same quantity? This is the question statisticians have asked and come up with reasonable ideas. One particular suggestion to follow here is the consolidation by conflation [^4] Throughout I will refer to this as “consolidation be conflation” or simply “consolidation”. due to T. P. Hill and J. Miller (Hill, 2011; Hill & Miller, 2011). In this approach the consolidated probability distribution of a variable $x$ is defined by
@@ 2.3 Binary Sample
This study requires a sample of pure binaries that are free of hidden close components. I apply a multi-step screening process to ensure that no (or extremely few) kinematic contaminated systems survive. I start with the initial selection criteria that were used by Chae (2024a, b). The initial selection is based on the requirement that binaries in a bin of sufficiently high accelerations must agree with the expectation of Newtonian dynamics in a statistical sense, specifically in terms of the median or average properties. Such a selection (Chae, 2024a, b) can be achieved by requiring sufficiently good precision of proper motions (PMs), parallaxes, and radial velocities (RVs) for both components and dynamical consistency between them. As already acknowledged in Chae (2024a, b), there can be (a small number of) individual exceptions that elude such a statistical selection. [^6] Minor bumps in the high-velocity tails of normalized projected velocity distributions in Figures 15 - 17 of Chae (2024c) may indicate such exceptions.
@@ 3.1 Overview and inspection of the individual Bayesian results
The 652 wide binaries in the statistical sample defined in Section 2.3 are individually modeled in the same manner as mock Newtonian binaries are modeled in Section 2.2. Each binary is modeled with both the thermal and the flat priors on eccentricity. Here, I examine the properties of the Bayesian modeling outputs before inferring gravity based on them in the following subsections.
@@
In all cases the individual PDFs of $\Gamma$ are much broader than the difference $\approx 0.07$ between Newton and MOND, and thus any specific result strongly inconsistent (e.g. $>3\sigma$) with the range $0\leq\Gamma\leq 0.07$ set by Newton and MOND may signal kinematic contaminants of hidden low-mass objects (e.g. faint stars or very massive Jovian planets) close to either or both of the stars. Although the statistical sample was selected through a rigorous screening process, rare exceptions may be present. As a specific example, consider an unobserved companion star with $0.1M_{\odot}$ orbiting a host star with $1.0M_{\odot}$ at a separation of 10 au. For the sake of simplicity, assume a circular orbit. Then, the relative velocity is $v\approx 9900$ m s-1 according to Equation (22) and the velocity perturbation experienced by the host star is $\approx 9900\times 0.1/1.1=900$ m s-1, which is significant. When the measured relative velocity between the two main stars is boosted by such a perturbation, the Bayesian modeling may return a PDF of $\Gamma$ that excludes the range $0\leq\Gamma\leq 0.07$ with a high significance. This is confirmed with a mock sample including hierarchical systems.
Figure 15: Same as Figure 14 but for wide binaries with suspicious kinematic contaminants. The posterior PDFs of $\Gamma$ for these systems are unlike all the other in that one of them alone would strongly exclude Newtonian gravity. But, more reasonable possibility is that they suffer from kinematic contamination. All eight of such cases are listed in Table 2.
Table 2: Wide binaries with suspected kinematic contaminants as inferred by Bayesian modeling
@@
Note. (1) The quoted values of $e$ and $\Gamma$ are from the posterior PDFs as examples are given in Figure 15. If these were pure wide binaries without hidden kinematic contaminants, one of these systems alone would rule out Newtonian gravity. (2) Most stars have normal ruwe values; ${\tt ruwe}<1.3$ are satisfied by all except for three stars, which have $1.3\lesssim{\tt ruwe}\lesssim 1.5$ that are still not so large. This indicates that kinematically abnormal systems not necessarily have large ruwe values.
@@
Any kinematic perturbers may cause outliers only in the positive direction of $\Gamma$. This is indeed the case for the Bayesian results of the statistical sample. To be concrete, I define the lower ($\sigma_{\rm{low}}$) and upper ($\sigma_{\rm{up}}$) uncertainties of $\Gamma$ with the 16th and 84th percentiles in its PDF. Then, $\Gamma_{\rm{med}}+3\sigma_{\rm{up}}>0$ is satisfied in all cases, but $\Gamma_{\rm{med}}-3\sigma_{\rm{down}}<0.02$ (here $0.02$ is used to allow for a median effect of RV uncertainties in a transition regime between the Newtonian and the MOND regimes) is not satisfied by eight wide binaries. Two examples of these outliers can be found in Figure 15. Clearly, any of these results alone would not only rule out the range $0\leq\Gamma\leq 0.07$, but would also be inconsistent with the PDFs of many other wide binaries. The properties of the eight wide binaries with suspected kinematic contaminants are summarized in Table 2. These systems will not be included in the inferring of consolidated gravity in the following subsections. In doing so, it is guaranteed that all individual binaries are consistent with Newtonian gravity within $\approx 3\sigma$ and the consolidated gravity is not dominated by a few radically different ones, but determined by the collective trend of the majority.
@@
Figure 16: Bayesian inferred eccentricities for the statistical sample from this work ($e$) are compared with those ($e_{\rm{H2022}}$) from Hwang et al. (2022). The upper row shows the eccentricities inferred with the thermal prior probability density $p(e)=2e$ while the lower row shows those with the flat prior. In the former case, there is an excellent match between this work and Hwang et al. (2022) as the middle column compares the median values ($e_{\rm{med}}$) from this work with the maximum likelihood estimate (MLE) values ($e_{\rm{MLE,H2022}}$) from Hwang et al. (2022). The right column compares “$1\sigma$” full widths between this work and Hwang et al. (2022). Because this work uses the additional information of radial velocities, the widths from this work are overall narrower.
@@
Properties of the inferred orbital and geometric parameters are of interest not only to check whether they are reasonable but also in its own right. Eccentricities of wide binaries are of particular interest. For all wide binaries in the El-Badry et al. (2021) catalog from which the statistical sample was drawn, Hwang et al. (2022) have inferred eccentricities with an independent Bayesian method using only the sky-projected relative displacements and relative velocities (without RVs). Figure 16 compares the eccentricities from this work with those from Hwang et al. (2022). The eccentricities with the thermal prior agree excellently with the Hwang et al. (2022) values, while those with the flat prior show some discrepancy. This is not surprising given that the statistical properties of observationally inferred eccentricities for wide binaries (see Figure 24 of Chae (2023a)) are much more consistent with the median eccentricity ($\sqrt{0.5}\approx 0.71$) of the thermal distribution than that ($0.5$) of the flat distribution. Moreover, as will be shown in Section 3.2 below, a test with the Newtonian control sample indicates that the flat prior can cause some bias in the Bayesian inference of gravity when RVs are not sufficiently precise. The right column of Figure 16 compares the widths of the PDFs between this work and Hwang et al. (2022). Because this work uses an additional constraint from RVs, it is expected that the widths from this work will be narrower. Indeed, this is the case. In particular, the results with the thermal prior exhibit narrower widths more clearly.

### v3

@@ Abstract
Isolated wide binary stars provide natural laboratories to directly probe gravity for Newtonian acceleration $g_{\rm{N}}\lesssim 10^{-9}$ m s-2. Recent statistical analyses of wide binaries have been performed only with sky-projected relative velocities $v_{p}$ in the pairs. A new method of Bayesian orbit modeling exploiting three relative velocity components including the radial (line-of-sight) component $v_{r}$ is developed to measure a gravitational anomaly parameter $\Gamma\equiv\log_{10}\sqrt{\gamma_{g}}\equiv\log_{10}\sqrt{G_{\rm{eff}}/G_{\rm{N}}}$ where $G_{\rm{eff}}$ is the effective gravitational constant for pseudo-Newtonian elliptical orbits, while $G_{\rm{N}}$ is Newton’s constant. The method infers individual probability distributions of $\Gamma$ and then combines the independent distributions to obtain a consolidated distribution in a specific range of $g_{\rm{N}}$. Here the method is described and applied to a sample of 312 wide binaries in a broad dynamic range $10^{-11.0}\lesssim g_{\rm{N}}\lesssim 10^{-6.7}$ m s-2 with $v_{r}$ uncertainties in the range $380168<\sigma_{v_{r}}<$ 380 m s-1 selected from the Gaia DR3 database. The following results are obtained: $\Gamma=0.000\pm 0.011$ or $\gamma_{g}=1.00\pm 0.05$ ($N_{\rm{binary}}=125$) for a high acceleration regime ($10^{-7.9}\lesssim g_{\rm{N}}\lesssim 10^{-6.7}$ m s-2) agreeing well with Newton, but $\Gamma=0.085\pm 0.040$ or $\gamma_{g}=1.48_{-0.23}^{+0.33}$ (35) for a MOND regime ($10^{-11.0}\lesssim g_{\rm{N}}\lesssim 10^{-9.5}$ m s-2) and $\Gamma=0.063\pm 0.015$ or $\gamma_{g}=1.34_{-0.08}^{+0.10}$ (111) for a MOND+transition regime ($10^{-11.0}\lesssim g_{\rm{N}}\lesssim 10^{-8.5}$ m s-2). These results show that gravitational anomaly is evident for $g_{\rm{N}}\lesssim 10^{-9}$ m s-2 and $\Gamma$ in the MOND regime ($\lesssim 10^{-9.5}$ m s-2) agrees with the first-tier prediction ($\Gamma\approx 0.07$ or $\gamma_{g}\approx 1.4$) of MOND gravity theories.
@@ 1 Introduction
The nature of gravity in the low acceleration limit of the nonrelativistic regime is of great importance because any deviation from standard gravity of Newton and Einstein represented by Poisson’s equation will have profound implications for the concept of dark matter, dynamics of astrophysical systems, cosmological theories, and fundamental theories of physics. For the past decade or so, wide binary stars have been recognized (Hernandez et al., 2012; Pittordis & Sutherland, 2018; Banik & Zhao, 2018; Hernandez et al., 2022) as potentially decisive testbeds of gravity at low acceleration because a significant number of wide binaries can be found in the solar neighborhoods, permitting relatively direct observational studies and their internal dynamics is free from the effects of hypothetical dark matter.
Several statistical analyses of wide binary internal motions carried out during 2023 - 2024 based on Gaia data release 3 (DR3; Vallenari et al., 2023), by Chae (Chae, 2023a, 2024a, 2024c) and independently by Hernandez et al. (Hernandez, 2023; Hernandez et al., 2024a; Hernandez & Kroupa, 2025), indicate that standard gravity is not only broken at low internal Newtonian acceleration ($g_{\rm{N}}\lesssim 10^{-9}$ m s-2) but gravity is boosted by approximately 40% at $g_{\rm{N}}\lesssim 10^{-10}$ m s-2 in agreement with the generic prediction of modified gravity theories (Bekenstein & Milgrom, 1984; Milgrom, 2010) in the paradigm of modified Newtonian dynamics (MOND; Milgrom, 1983). Although divergent results (Banik et al., 2024) were published during the same period, the recent critical review by Hernandez et al. (2024b) (see also Chae 2024c for more details) addresses the origin of the divergent results, in particular the lack of the self-calibration procedure for the fraction of hierarchical systems in an impure binary sample (i.e. including hierarchical systems).
@@
This paper is structured as follows. In Section 2, I describe the methodology and data. In detail, I first develop a Bayesian methodology to infer probability density functions (PDFs) of gravity from individual binaries (Section 2.1), and then describe the method of statistically combining individual PDFs to obtain a consolidated probability distribution by maximizing information (Section 2.2). Section 2.3 describes the selection of a pure binary sample from the Gaia DR3 database that will be used for individual Bayesian modeling. The selection of the sample will be more stringent than previous samples that were used mainly for median statistics in bins of accelerations (or equivalent variables) by Chae (2024a, c). In Section 3, I describe the results of Bayesian modeling based on the Gaia sample. In Section 3.1, I first examine the individual Bayesian results and the overall properties of the results. In Section 3.2, I present the consolidated probability distributions of gravity with wide binaries in a Newtonian (relatively strong internal acceleration) regime. In Section 3.3, I present the main results of this work, i.e. the consolidated probability distributions of gravity in low-acceleration regimes. In Section 4, I discuss the meaning of the results and future wide binary gravity research. In Section 5, I present the current conclusion and prospects with future data on radial velocities. In Appendix A, a general transformation with the Euler angles between the orbital plane and the observer’s frame is described as it is needed to produce mock samples. In Appendix B, perspective effects in proper motions and radial velocities of wide binary stars are investigated to explore any possible biases in the low-acceleration regime data.
For a quick grasp of this rather lengthy paper, the reader may read Sections 2.1, 2.2, 3.2, 3.3, 4.1, and 5. The wide binary samples and the main Bayesian modeling results are available on Zenodo under an open-source Creative Commons Attribution license: https://doi.org/10.5281/zenodo.14900456 (catalog 10.5281/zenodo.14900456). The key Python codes are also available (Chae, 2025).
@@ 2.1 Bayesian 3D modeling
The above equations (Equations (5) and (7)) for the relative velocity components based on the observed PMs and RVs do not include corrections for the “perspective effects” (Shaya & Olling, 2011) that are apparent relative velocities between the pair in the observer’s spherical coordinates due to the motion of the barycenter of the system in the 3D space. For wide binaries in the solar neighborhood, perspective effects can be relevant for $s>0.1$ pc (i.e. 20 kau) (El-Badry, 2019).
Almost all binaries considered in this work have $s<20$ kau (99% satisfy $s<15.5$ kau) meaning that perspective effects are minimal, and thus their corrections are not warranted when RV uncertainties are large. Moreover, for an accurate estimate of the perspective effects, the difference in distances of the two stars is needed. Unfortunately, observed distance uncertainties are (usually much) larger than $s$ and one has to assume zero (or reasonably small) difference in the distances to obtain a first-order estimate of the perspective effects. The default choice of the present work will be to ignore any perspective effects. However, I will also consider the case of correcting for first-order perspective effects estimated by a numerical procedure described in Appendix B.
@@
which is related to the inclination angle $i$ by $\tau_{\rm{obs}}=\tan i$. Finally, the magnitude of the relative physical velocity in the 3D space is then given by
@@
In Equations (10), (12), and (13), $e$ is the eccentricity, $i$ is an inclination parameter, [^1] I note that this parameter is not the same as the angle between the orbital plane and the plane of the sky. $\phi_{0}$ is the longitude of the periastron, and $\phi$ is the phase angle given by
$\phi=\tan^{-1}((\Delta y^{\prime}/\Delta x^{\prime})/\cos i),$ (14)
which is essentially a function of only $i$ in the observer’s frame because $\Delta x^{\prime}$ and $\Delta y^{\prime}$ are precisely measured. In Equation (10), $\gamma_{g}$ is the gravity boost factor defined in Equation (1). In Equation (11), $M_{\rm{tot}}$ is the total mass of the binary system to be constrained, while $M_{\rm{tot,obs}}$ is the observational estimate (see Section 2.3). The set of free parameters for each binary is thus given by $\mathbf{\Theta}=\{e,\phi_{0},i,\log_{10}(f_{M}),\Gamma\}$.
@@
For eccentricity $e$ I consider the range $0.9990.001<e<$ 0.999 with a prior probability distribution. For the prior, I consider three choices: a flat prior, a “thermal” probability density $p(e)=2e$, or a separation($s$)-dependent probability density $p(e)=(1+\alpha)e^{\alpha}$ with $\alpha=\alpha(s)$ from Hwang et al. (2022) (hereafter referred to as the “H2022” prior) as described by Equation (18) of Chae (2024a). If the data can determine the orbit well, the flat prior may be sufficient. However, with snapshot data, $e$ is ill-constrained in particular for relatively large uncertainties of $v_{r}$ as in many cases based on the current Gaia DR3 database. When $e$ is ill-constrained (or essentially unconstrained in some cases), the prior matters because the flat prior would be equivalent to assuming that the median eccentricity is $0.5$ while the thermal prior has a median of $\sqrt{0.5}(\approx 0.71)$ more consistent with the observed statistical properties. The H2022 prior has $\alpha\approx 1.3$ for $s>1$ kau(kilo astronomical unit) implying a median of $\approx 0.74$ that differs from the thermal case by just $0.03$, indicating that two priors would make little difference. Moreover, because individual data themselves have some constraining power in this work, an eccentricity probability distribution plays a weaker role as a prior than a fixed input as in previous studies. For this reason, the thermal prior will be the default choice.
@@
where $\phi$ is given by Equation (14). The parameter $\phi_{0}$ is allowed to vary freely within the range $(-360^{\circ},360^{\circ})$.
For $f_{M}$, I assume a Gaussian probability density for $\log_{10}(f_{M})$ with $(\mu,\sigma)=(0,0.021)$ allowing a 5% scatter of $M_{\rm{tot}}$, which covers uncertainties of mass estimates (see Section 2.3). For $i$ and $\Gamma$, I assume flat priors in the ranges: $-90^{\circ}<i<90^{\circ}$ and $-1<\Gamma<1$.
@@
Given that the averaged PDF is not satisfactory, what would be the best way to extract the maximal information from the set of individual PDFs, or how to combine independent measurements of the same quantity? This is the question statisticians have asked and come up with reasonable ideas. One particular suggestion to follow here is the consolidation by conflation [^4] Throughout I will refer to this as “consolidation by conflation” or simply “consolidation”. due to T. P. Hill and J. Miller (Hill, 2011; Hill & Miller, 2011). In this approach the consolidated probability distribution of a variable $x$ is defined by
@@ 2.3 Binary Sample
This study requires a sample of pure binaries that are free of hidden close components. Selecting and using pure binaries was originally considered by Hernandez et al. (2022) and Hernandez (2023) in the context of statistical tests of gravity with $v_{p}$. Here I apply a multi-step screening process to ensure that no (or extremely few) kinematic contaminated systems survive. I start with the initial selection criteria that were used by Chae (2024a, b). The initial selection is based on the requirement that binaries in a bin of sufficiently high accelerations must agree with the expectation of Newtonian dynamics in a statistical sense, specifically in terms of the median or average properties. Such a selection (Chae, 2024a, b) can be achieved by requiring sufficiently good precision of proper motions (PMs), parallaxes, and radial velocities (RVs) for both components and dynamical consistency between them. Note that availability of Gaia measured RVs (with a reasonable precision) particularly helps removing systems with hidden close companions because it implies a good single-star solution for the spectroscopic observations. However, as already acknowledged in Chae (2024a, b), there can be (a small number of) individual exceptions that elude such a statistical selection. [^6] Minor bumps in the high-velocity tails of normalized projected velocity distributions in Figures 15 - 17 of Chae (2024c) may indicate such exceptions.
@@
For the binary stars of the selected samples stellar masses are estimated using the standard $M_{G}$-mass relation obtained by Chae (2023a) (the first choice in Table 1 of that work) where $M_{G}$ is the $G$-band absolute magnitude corrected for dust extinction as described in Chae (2023a). The left panel of Figure 13 shows the distribution of the Chae (2023a) masses of 1304 stars in the statistical sample. For these masses that are mainly $\gtrsim 0.5M_{\odot}$, the uncertainties of 5% may well encompass alternative masses as shown in Figure 7 of Chae (2023a). For alternative masses, it is relevant to consider those provided by the Gaia DR3 Final Luminosity Age Mass Estimator (FLAME). FLAME masses are independent and individual results for selective stars from stellar evolution modeling based on a host of Gaia photometric, spectroscopic, and astrometric data. FLAME masses are available for only 859 out of the 1304 stars. The right panel of Figure 13 shows the ratio of the FLAME mass to the Chae (2023a) mass for the subset of stars. The mean and standard deviation of the ratios are $0.97$ and $0.05$ indicating that two masses are indeed consistent with each other up to about 5% scatter. Because FLAME masses are not available for all stars, masses in this work refer to the Chae (2023a) values unless stated otherwise.
Figure 13: The left panel shows the distribution of the Chae (2023a) masses ($M_{\rm{C2023}}$) for all stars of the statistical sample. The right panel compares the Gaia FLAME masses ($M_{\rm{FLAME}}$) with $M_{\rm{C2023}}$ for a subset of stars that have $M_{\rm{FLAME}}$.
@@ 3.1 Overview and inspection of the individual Bayesian results
The 652 wide binaries in the statistical sample defined in Section 2.3 are individually modeled in the same manner as mock Newtonian binaries are modeled in Section 2.2. Each binary is modeled with the thermal, the H2022, and the flat priors on eccentricity. As will be shown below, the results with the thermal and H2022 priors are very similar while those with the flat prior are mildly different. Here, I examine the properties of the Bayesian modeling outputs before inferring gravity based on them in the following subsections. In presenting results below, the thermal prior is implicit unless specified otherwise.
@@
Figure 16: Same as Figure 15 but for wide binaries with abnormal PDFs of $\Gamma$. The posterior PDFs of $\Gamma$ for these systems are unlike all the other in that one of them alone would strongly exclude Newtonian gravity. But, more reasonable possibility is that they suffer from kinematic contamination. All eight of such cases are listed in Table 2.
In all cases the individual PDFs of $\Gamma$ are much broader than the difference $\approx 0.07$ between Newton and MOND, and thus any specific result strongly inconsistent (e.g. $>3\sigma$) with the range $0\leq\Gamma\leq 0.07$ set by Newton and MOND may signal kinematic contaminants of hidden low-mass objects (e.g. faint stars or very massive Jovian planets) close to either or both of the stars, or some unknown issues with the RV data999 See Katz et al. (2023) for a general description of the Gaia DR3 RV data.. Although the statistical sample was selected through a rigorous screening process, rare exceptions may be present. As a specific example, consider an unobserved companion star with $0.1M_{\odot}$ orbiting a host star with $1.0M_{\odot}$ at a separation of 10 au. For the sake of simplicity, assume a circular orbit. Then, the relative velocity is $v\approx 9900$ m s-1 according to Equation (23) and the velocity perturbation experienced by the host star is $\approx 9900\times 0.1/1.1=900$ m s-1, which is significant. When the measured relative velocity between the two main stars is boosted by such a perturbation, the Bayesian modeling may return a PDF of $\Gamma$ that excludes the range $0\leq\Gamma\leq 0.07$ with a high significance. This is confirmed with a mock sample including hierarchical systems.
Table 2: Wide binaries with abnormal PDFs of $\Gamma$ as inferred by Bayesian modeling
@@
Note. (1) The quoted values of $e$ and $\Gamma$ are from the posterior PDFs as examples are given in Figure 16. If these were pure wide binaries without hidden kinematic contaminants and the RV values and uncertainties were reliable, one of these systems alone would rule out Newtonian gravity. (2) Most stars have normal ruwe values; ${\tt ruwe}<1.3$ are satisfied by all except for three stars, which have $1.3\lesssim{\tt ruwe}\lesssim 1.5$ that are still not so large. This indicates that kinematically abnormal systems based on Gaia data not necessarily have large ruwe values.
@@
Any kinematic perturbers may cause outliers only in the positive direction of $\Gamma$. This is indeed the case for the Bayesian results of the statistical sample. To be concrete, I define the lower ($\sigma_{\rm{low}}$) and upper ($\sigma_{\rm{up}}$) uncertainties of $\Gamma$ with the 16th and 84th percentiles in its PDF. Then, $\Gamma_{\rm{med}}+3\sigma_{\rm{up}}>0$ is satisfied in all cases, but $\Gamma_{\rm{med}}-3\sigma_{\rm{down}}<0.02$ (here $0.02$ is used to allow for a median effect of RV uncertainties in a transition regime between the Newtonian and the MOND regimes) is not satisfied by eight wide binaries. Two examples of these outliers can be found in Figure 16. Clearly, any of these results alone would not only rule out the range $0\leq\Gamma\leq 0.07$, but would also be inconsistent with the PDFs of many other wide binaries. The properties of the eight wide binaries with suspected kinematic contaminants or abnormal RV data are summarized in Table 2. These systems will not be included in the inferring of consolidated gravity in the following subsections. In doing so, it is guaranteed that all individual binaries are consistent with Newtonian gravity within $\approx 3\sigma$ and the consolidated gravity is not dominated by a few radically different ones, but determined by the collective trend of the majority. See, however, Section 4.1 for discussions on alternative interpretations.
@@
Figure 17: Bayesian inferred eccentricities for the statistical sample from this work ($e$) are compared with those ($e_{\rm{H2022}}$) from Hwang et al. (2022). The upper, middle, and lower rows show respectively the eccentricities inferred with the thermal, the H2022, and the flat priors. In the upper and middle rows, there is an excellent match between this work and Hwang et al. (2022) as the middle column compares the median values ($e_{\rm{med}}$) from this work with the maximum likelihood estimate (MLE) values ($e_{\rm{MLE,H2022}}$) from Hwang et al. (2022). The right column compares “$1\sigma$” full widths between this work and Hwang et al. (2022). Because this work uses the additional information of radial velocities, the widths from this work are overall narrower.
@@
Properties of the inferred orbital and geometric parameters are of interest not only to check whether they are reasonable but also in its own right. Eccentricities of wide binaries are of particular interest. For all wide binaries in the El-Badry et al. (2021) catalog from which the statistical sample was drawn, Hwang et al. (2022) have inferred eccentricities with an independent Bayesian method using only the sky-projected relative displacements and relative velocities (without RVs). Figure 17 compares the eccentricities from this work with those from Hwang et al. (2022). The eccentricities with the thermal prior agree excellently with the Hwang et al. (2022) values as those with the H2022 prior do. However, those with the flat prior show some discrepancy. This is not surprising given that the statistical properties of observationally inferred eccentricities for wide binaries (see Figure 24 of Chae (2023a)) are much more consistent with the thermal or H2022 distribution than the flat distribution. Moreover, as will be shown in Section 3.2 below, a test with the Newtonian control sample indicates that the flat prior can cause some bias in the Bayesian inference of gravity when RVs are not sufficiently precise. The right column of Figure 17 compares the widths of the PDFs between this work and Hwang et al. (2022). Because this work uses an additional constraint from RVs, it is expected that the widths from this work will be narrower. Indeed, this is the case. In particular, the results with the thermal or H2022 prior exhibit narrower widths more clearly.